OpenStudy (anonymous):

what do 3arasi mean in arabic ???

OpenStudy (anonymous):

3arasi ? in which text you find this?

OpenStudy (aroub):

lol

OpenStudy (aroub):

it cant be translated to English..

OpenStudy (anonymous):

it can be so what does?????????

OpenStudy (anonymous):

it mean

OpenStudy (anonymous):

i think it means donkey !!!

OpenStudy (anonymous):

or bio nerd

OpenStudy (anonymous):

or even stalker

OpenStudy (anonymous):

!!!!!!!!!!!!!!!!!! ;)

OpenStudy (anonymous):

\[f''_-(r)=2a\] \[f''_+(r)=-2\cos(2r)\] so we need \[2a=-2\] \[a=-1\]

OpenStudy (anonymous):

so far we have \[f_-(r)=-r^2+br+3\]

OpenStudy (anonymous):

now how to get b...

OpenStudy (anonymous):

oh i see \[f'_-(r)=-2r+b\] and \[f'_+(r)=f'_+(0)=4\sin(r)\cos(r)-6\cos(r)\sin(r)\]

OpenStudy (anonymous):

typo

OpenStudy (aroub):

ok, ill translated but it sounds weird in english.. 3arasi=on my head dont ask!

OpenStudy (anonymous):

on ur head !!!!!

OpenStudy (anonymous):

but in any case \[b=0\]

OpenStudy (aroub):

or satellites answer..

OpenStudy (anonymous):

which if i was not a bone head we could have got right away

OpenStudy (anonymous):

repost and i will write it step by step

OpenStudy (anonymous):

ok aroub is right this is one of the rabic accent

OpenStudy (anonymous):

but why you asked it here in this site?

OpenStudy (anonymous):

bye and fe aman illah !!!!

OpenStudy (lalaly):

lol hashir ur crazy

OpenStudy (anonymous):

o knw i m !!!!

OpenStudy (anonymous):

o=I

OpenStudy (lalaly):

hehehe xP

OpenStudy (aroub):

hehehe fee aman illah ...